OpenStudy (anonymous):

x^2+√3x-1+ how do I solve this?

OpenStudy (anonymous):

\[x^2+\sqrt{3}x-1=0\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=\sqrt{3},c=-1\]

OpenStudy (anonymous):

get \[x=\frac{-\sqrt{3}\pm\sqrt{3-4\times 1\times -1}}{2}\]

OpenStudy (anonymous):

or \[x=\frac{-\sqrt{3}\pm\sqrt{7}}{2}\]

OpenStudy (anonymous):

so the square root under the square root cancels out?

OpenStudy (anonymous):

? nothing cancels

OpenStudy (anonymous):

maybe you mean when you compute \[b^2\] with \[b=\sqrt{3}\] you get \[\sqrt{3}^2=3\] which is right

OpenStudy (anonymous):

yeah duh, I freaked out for no reason. Thanks!!!

OpenStudy (anonymous):

yw